# An $\ell^{1}$-Plug-and-Play Approach for MPI Using a Zero Shot Denoiser with Evaluation on the 3D Open MPI Dataset

### v2

@@ 3.2 Methods for Comparison
The PP-MPI Algorithm with Trained Denoiser
@@
where $\alpha_{i}$ is the parameter weighting the i-th regularization function $g_{i}$, the data fidelity term for given data $f$ and is imposed as a constraint; $\epsilon$ is the upper bound on the noise level. The constrained minimization problem in Equation (19) can be solved using an ADMM splitting scheme that yields the following algorithm:
$\begin{split}u^{n+1}&=\left(\mathbb{I}+A^{*}A\right)^{-1}\left(A^{*}(z_{0}^{n}+d_{0}^{n})+z_{1}^{n}+d_{1}^{n}\right)\\ z_{1}^{n+1}&=\mathrm{Denoiser}(u^{n+1}-d_{1}^{n})\\ d_{1}^{n+1}&=d_{1}^{n}+z_{1}^{n+1}-u^{n+1}\\ z_{0}^{n+1}&=\Psi_{f,\epsilon}(Au^{n+1}-d_{0}^{n})\\ d_{0}^{n+1}&=d_{0}^{n}+z_{0}^{n+1}-Au^{n+1}\\ \end{split}$ (20)
where $\Psi_{f,\epsilon}$ is the projection function onto the ball of radius $\epsilon$ centered at $f$, taking care of the data fidelity constraint. The denoiser trained by the authors [4] on MPI-friendly data will be called PP-MPI-Denoiser in what follows.
3.3 Experiment 1: Comparison on the OpenMPI Dataset
@@
Parameter Validation
We display results for the proposed ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP approach, the DIP as well as the Tikhonov regularization method and the PP-MPI approach.
For the comparison method of Tikhonov regularization, we have chosen its parameter $\lambda_{\mathrm{Tik}}$ optimally in the sense that we have validated it performing a grid search and selecting that value for which the average $\mathrm{PSNR}$ is maximal. The grid search proceeds as follows: we first perform reconstruction with parameters $10^{j}$ for $j=-6,-5,\dots,0,1,\cdots,18$ and consider $j^{\ast}$ the exponent that maximizes the PSNR; then we consider the values $k\cdot 10^{j^{\ast}-1}$ and $k\cdot 10^{j^{\ast}}$ with $k=1,\dots,9$ and again select the one maximizing the average PSNR.
For the fine-tuning of the DIP comparison method, we have followed the approach proposed in [14]: we let the DIP run for 20000 iterations with three different gradient descent step sizes $\alpha_{i}=10^{-i}$, $i=3,4,5$ on the validation dataset and we select the parameter $\alpha_{i}$ and the iteration $s$ for which the average $\mathrm{PSNR}$ is maximized. (We refer to Section 3.2 and [14] for details on the implementation.)
For the PP-MPI in  [4] we have validated the number of iterations and the variable $\epsilon$ in Equation (19). Consistently with [4], we consider $f=Au_{\mathrm{GT}}+\varepsilon\eta$ for $\eta\sim\mathcal{N}(0,\mathbb{I})$ with $\varepsilon=\sqrt{\lVert Au_{\mathrm{GT}}\rVert^{2}/(\mathrm{SNR}\cdot\lVert\eta\rVert^{2})}$, and $\mathrm{SNR}_{\mathrm{db}}=10\log_{10}(\mathrm{SNR})$. We performed 3000 total iterations with $\epsilon$ with $\mathrm{SNR}_{db}$ ranging from 20 to 30 with step size 1.
For both the ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP methods we have validated the parameter $\mu_{0}$ and the optimal number of iterations $n_{\mathrm{it}}$ again considering the optimal parameters that maximize the average $\mathrm{PSNR}$. For $\mu_{0}$ we have performed the same grid search as for the Tikhonov parameter.
Comparison Methods Proposed ZeroShot-PnP
Tikhonov DIP PP-MPI [4] Without $\ell^{1}$-Prior With $\ell^{1}$-Prior
Param. $\lambda_{\mathrm{Tik.}}=5\cdot 10^{5}$ $\alpha=10^{-3}$ $\mathrm{SNR}_{db}=25$ $\mu_{0}=2\cdot 10^{5}$ $\mu_{0}=3\cdot 10^{5}$
$s=14000$ $n_{\mathrm{it}}=2170$ $n_{\mathrm{it}}=7$ $n_{\mathrm{it}}=5$
PSNR $25.28\pm 3.28$ $29.45\pm 3.07$ $26.10\pm 3.95$ $30.45\pm 5.04$ $29.77\pm 5.82$
SSIM $0.657\pm 0.140$ $0.646\pm 0.215$ $0.524\pm 0.195$ $0.813\pm 0.140$ $0.794\pm 0.187$
@@
Table 1: Validation results on the hybrid dataset in Section 2.5. We observe that the proposed ZeroShot-PnP methods achieve the highest values w.r.t. both PSNR and SSIM on the hybrid datset.
An overview of the validated parameters is displayed in Table 1. We observe that on the validation dataset both the ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP proposed in this paper achieve higher $\mathrm{PSNR}$ and $\mathrm{SSIM}$ values.
(a) Shape CAD reference.
(b) Shape Tikhonov.
(c) Shape DIP.
(d) Shape PP-MPI.
(e) Shape ZeroShot-PnP.
(f) Shape ZeroShot-$\ell^{1}$-PnP.
@@
(a) Res. CAD reference.
(b) Res. Tikhonov.
(c) Res. DIP.
(d) Res. PP-MPI.
(e) Res. ZeroShot-PnP .
(f) Res. ZeroShot-$\ell^{1}$-PnP.
@@
(a) Conc. CAD reference.
(b) Conc. Tikhonov.
(c) Conc. DIP.
(d) Conc. PP-MPI.
(e) Conc. ZeroShot-PnP.
(f) Conc. ZeroShot-$\ell^{1}$-PnP.
@@
Results
Selected slices of the reconstructions of the three phantoms are displayed in Figures 12 and 3. The reconstructions where performed with the validated parameters in Table 1.
@@ Results
Tikhonov DIP PP-MPI [4] with $\ell^{1}$-Prior without $\ell^{1}$-Prior
Shape $\mathrm{PSNR}_{\mathrm{max}}$ 22.64 28.96 21.14 31.84 31.87
Phatom $\mathrm{SSIM}_{\mathrm{max}}$ 0.489 0.871 0.769 0.954 0.954
Resolution $\mathrm{PSNR}_{\mathrm{max}}$ 30.64 32.89 30.09 32.1 32.09
Phantom $\mathrm{SSIM}_{\mathrm{max}}$ 0.593 0.737 0.519 0.731 0.730
Concentr. $\mathrm{PSNR}_{\mathrm{max}}$ 36.14 37.78 36.09 37.54 37.45
Phatom $\mathrm{SSIM}_{\mathrm{max}}$ 0.518 0.441 0.535 0.578 0.577
Table 2: Results obtained with the ZeroShot-$\ell^{1}$-PnP, the Tikhonov, the DIP and the PP-MPI methods on the 3D Open MPI dataset. We display the $\mathrm{PSNR}_{\mathrm{max}}$ and $\mathrm{SSIM}_{\mathrm{max}}$ scores for the reconstructions obtained using the validated parameter in Table 1.
An overview of the results obtained in terms of $\mathrm{PSNR}_{\mathrm{max}}$ and $\mathrm{SSIM}_{\mathrm{max}}$ are displayed in Tab. 2.
We can observe that the ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP methods achieve similar results, and in some instances the additional $\ell^{1}$-prior can increase the PSNR score (cf. shape phantom). In any case, the reconstruction quality according to both PSNR and SSIM of ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP are higher than the baseline Tikhonov method and the previous PP-MPI approach. The reconstruction quality of the two methods proposed in this paper for MPI reconstruction are comparable in quality with the values obtained for the DIP. However, the DIP takes considerably more time to perform a reconstruction and involves a random initial state $z$ on which the quality of the reconstruction depends. Moreover, comparing the DIP reconstruction in Figure 3(c) with the CAD reference 3(a) and the ZeroShot-$\ell^{1}$-PnP reconstruction 3(f), it is possible that even if the PSNR value is the highest, some features of the phantom may not be reconstructed by the DIP.
@@
Because the $19\times 19\times 19$ resolution of the reconstructions, the visual quality assessment of the differences between reconstructions can be difficult. For a better understanding of the differences we have plotted in Figure 4 a 1D line plot that compares the graph of the reconstructions and the reference CAD model. In Figure 4 we can see that the Tikhonov method smooths out the reference and takes negative values, whereas the the DIP and the PP-MPI tend to overshoot when compared to the ZeroShot-PnP reconstruction. The ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP lines are almost completely overlapping. We observe that the DIP can present artifacts in the reconstruction (cf. Figure 3(c) and 3(a)).
Runtimes
@@
Table 3: Wall clock runtime table of the examined methods. The times are computed by averaging 100 reconstructions on the Open MPI dataset. We observe the proposed ZeroShot-PnP and ZeroShot-$\ell^{1}$-PnP take the least amount of time for the reconstructions.
@@
3.4 Experiment 2: Study on the Differences with the PP-MPI Approach
In this experiment we study more closely the differences between the previous PP-MPI approach and the ZeroShot-PnP approaches proposed for MPI in this paper. In particular, we consider the PP-MPI algorithm with its denoiser trained on MPI-friendly data (PP-MPI-Denoiser) [4]; then, we consider the proposed Algorithm 1 first with the PP-MPI-Denoiser and then with the ZeroShot-Denoiser [70]. The objective here is to show that the employment of the zero-shot denoiser achieves competitive results when compared with the previous PPP-MPI-Denoiser. Moreover, this experiment shows that with the reconstruction parameter automatically selected as proposed in Section 2.4, it is possible to enhance the reconstruction results also if the PP-MPI-Denoiser is employed.
Preprocessing
@@ 3.4 Experiment 2: Study on the Differences with the PP-MPI Approach
Parameter Validation
We choose the parameters according to Section 2.4. The validated starting parameters $\mu_{0}$ as well as the resulting number of iterations are displayed in Table 4.
PP-MPI-Denoiser [4]
PP-MPI HQS-Splitting Eq.(12) $\ell^{1}$-HQS-Splitting Eq. (9)-(11)
Param. $\mathrm{SNR}_{db}=25$ $\mu_{0}=4\cdot 10^{11}$ $\mu_{0}=6\cdot 10^{16}$
Iteration $n_{\mathrm{it}}=2170$ $n_{\mathrm{it}}=9$ $n_{\mathrm{it}}=4$
PSNR $26.10\pm 3.95$ $27.75\pm 5.20$ $27.68\pm 5.15$
SSIM $0.524\pm 0.195$ $0.567\pm 0.224$ $0.644\pm 0.215$
Proposed Denoiser [70]
ZeroShot-PnP ZeroShot-$\ell^{1}$-PnP
Param. $\mu_{0}=2\cdot 10^{5}$ $\mu_{0}=3\cdot 10^{5}$
Iteration $n_{\mathrm{it}}=7$ $n_{\mathrm{it}}=5$
PSNR $30.45\pm 5.04$ $29.77\pm 5.82$
SSIM $0.813\pm 0.140$ $0.794\pm 0.187$
Table 4: Validation results on the hybrid dataset in Section 2.5. We observe that the employment of Algorithm 1 in combination with the PP-MPI-Denoiser leads to a lower number of iterations and to increased PSNR and SSIM scores when compared with the PP-MPI algorithm. However, the highest PSNR and SSIM score are obtained using the proposed ZeroShot-Denoiser (which has not been trained on MPI data).
Results
From the PSNR and SSIM values obtained on the hybrid dataset in Table 4 we can observe that both the HQS-Splitting (Equation (12)) and the $\ell^{1}$-HQS-Splitting of Equations (9)-(11) with the PP-MPI-Denoiser achieve higher metric scores than the original ADMM-based PP-MPI algorithm. Moreover, these improvements are achieved in less than 10 iterations in both cases. We observe that in the algorithm with the PP-MPI-Denoiser, the HQS-Splitting algorithm achieves the highest PSNR value whereas the $\ell^{1}$-HQS-Splitting achieves the highest SSIM score without compromising too much the PSNR: this means that there are cases in which one might prefer one method over the other depending on the objective. However, from the results in Table 4 it is also clear that the combination of the proposed splittings and the proposed denoiser [70], i.e., the ZeroShot-PnP and the ZeroShot-$\ell^{1}$-PnP approaches, yields the results with the highest PSNR and SSIM scores.
3.5 Experiment 3: Reconstructions with Decreasing Levels of Preprocessing for the ZeroShot-$\ell^{1}$-PnP
@@
ZeroShot PSNR 31.84 31.05 31.30
PnP SSIM 0.954 0.946 0.950
$n_{\mathrm{it}}$ 6 19 8
$\mu_{0}$ $2\cdot 10^{5}$ $6\cdot 10^{16}$ $10^{9}$
ZeroShot PSNR 31.87 31.14 29.8
$\ell^{1}$-PnP SSIM 0.954 0.947 0.956
$n_{\mathrm{it}}$ 4 3 74
@@
Table 5: The ZeroShot-$\ell^{1}$-PnP adn ZeroShot-PnP validated parameters and metric computed on the shape phantom with different levels of preprocessing.
Results

### v3

@@ 3.2 Methods for Comparison and Parameter Validation
The PP-MPI Algorithm with Trained Denoiser.
@@
where $u$ is the unknown distribution of particles, the $\alpha_{i}$ are the parameters weighting each i-th regularization function $g_{i}$, the data fidelity term for given scan data $f$ and system matrix $A$ is imposed as a constraint; $\epsilon$ is the upper bound on the noise level. The constrained minimization problem in Equation (23) can be solved using an ADMM splitting scheme that yields the following algorithm:
$\begin{split}u^{n+1}&=\left(\mathbb{I}+A^{T}A\right)^{-1}\left(A^{T}(z_{0}^{n}+d_{0}^{n})+z_{1}^{n}+d_{1}^{n}\right)\\ z_{1}^{n+1}&=\mathrm{Denoiser}(u^{n+1}-d_{1}^{n})\\ d_{1}^{n+1}&=d_{1}^{n}+z_{1}^{n+1}-u^{n+1}\\ z_{0}^{n+1}&=\Psi_{f,\epsilon}(Au^{n+1}-d_{0}^{n})\\ d_{0}^{n+1}&=d_{0}^{n}+z_{0}^{n+1}-Au^{n+1}\\ \end{split}$ (24)
where $\Psi_{f,\epsilon}$ is the projection function onto the ball of radius $\epsilon$ centered at $f$, taking care of the data fidelity constraint, and $z_{0}^{n}$, $z_{1}^{n}$, $d_{0}^{n}$, $d_{1}^{n}$ are auxiliary variables arising from the ADMM splitting that can be initialized as zero vectors at iteration $n=0$. The denoiser trained by the authors of [4] on MPI-friendly data will be called PP-MPI-Denoiser in what follows. For the PP-MPI algorithm we have used the algorithm and denoiser provided by the authors [4, 26] available at https://github.com/icon-lab/PP-MPI. We have validated the number of iterations and the variable $\epsilon$ in Equation (23) on the hybrid dataset. Consistently with [4], we consider $f=Au_{\mathrm{GT}}+\varepsilon\eta$ for $\eta\sim\mathcal{N}(0,\mathbb{I})$ with $\varepsilon=\sqrt{\lVert Au_{\mathrm{GT}}\rVert^{2}/(\mathrm{SNR}\cdot\lVert\eta\rVert^{2})}$, and $\mathrm{SNR}_{\mathrm{db}}=10\log_{10}(\mathrm{SNR})$. We performed 3000 total iterations with $\epsilon$ with $\mathrm{SNR}_{\mathrm{db}}$ ranging from 20 to 30 with step size 1.
Validation of the Proposed ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP Methods.
We have validated the parameter $\mu_{0}$ and the optimal number of iterations $n_{\mathrm{it}}$ considering the optimal parameters that maximize the average $\mathrm{PSNR}$ on the hybrid dataset. For $\mu_{0}$ we have applied the same grid search strategy as for the Tikhonov parameter above.
3.3 Experiment 1: Comparison on the Open MPI Dataset.
@@
Parameter Validation.
Comparison Methods
Tikhonov ART DIP PP-MPI [4]
Param. $\lambda_{\mathrm{Tik.}}=5\cdot 10^{5}$ $\lambda_{\mathrm{ART}}=2\cdot 10^{4}$ $\alpha=10^{-3}$ $\mathrm{SNR}_{\mathrm{db}}=$22
it=200 $s=14000$ $n_{\mathrm{it}}=$1606
PSNR $25.28\pm 3.28$ $27.00\pm 4.51$ $29.45\pm 3.07$ $28.13\pm 3.75$
SSIM $0.640\pm 0.147$ $0.482\pm 0.198$ $0.637\pm 0.248$ $0.408\pm 0.217$
Proposed ZeroShot-PnP
Without $\ell^{1}$-Prior With $\ell^{1}$-Prior
Param. $\mu_{0}=2\cdot 10^{5}$ $\mu_{0}=3\cdot 10^{5}$
$n_{\mathrm{it}}=7$ $n_{\mathrm{it}}=5$
PSNR $30.45\pm 5.04$ $29.77\pm 5.82$
SSIM $0.788\pm 0.141$ $0.766\pm 0.189$
@@
Table 1: Validation results on the hybrid dataset of Section 2.5. We observe that the proposed ZeroShot-PnP methods achieve the highest values w.r.t. both PSNR and SSIM on the hybrid datset.
(a) ZeroShot-PnP vs. Tikhonov.
(b) ZeroShot-PnP vs. ART.
(c) ZeroShot-PnP vs. DIP.
(d) ZeroShot-PnP vs. PPMPI.
(e) ZeroShot-$\ell^{1}$-PnP vs. Tikhonov.
(f) ZeroShot-$\ell^{1}$-PnP vs. ART.
(g) ZeroShot-$\ell^{1}$-PnP vs. DIP.
(h) ZeroShot-$\ell^{1}$-PnP vs. PPMPI.
Figure 1: Plot of the PSNR and SSIM distributions on the hybrid validation dataset representing the distribution of PSNR (in green and blue) and SSIM (in ocher and violet) in the ZeroShot-$\ell^{1}$-PnP (resp. ZeroShot-PnP) and the comparison methods. On the x-axis we display the ranges of the PSNR values (in $[0,40]$) and SSIM values (in $[0,1]$). the distributions have been estimated using Kernel Densities Estimations using Gaussian kernels.
As dataset for parameter validation we employ the hybrid dataset of Section 2.5. We display the found parameters for the proposed ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP approach, the DIP as well as the Tikhonov regularization method, the ART method and the PP-MPI approach. An overview of the validated parameters is displayed in Table 1. We observe that on the validation dataset both the ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP proposed in this paper achieve higher $\mathrm{PSNR}$ and $\mathrm{SSIM}$ values. In order to better understand the variation of the PSNR and SSIM values of the proposed ZeroShot-PnP and ZeroShot-$\ell^{1}$-PnP algorithms and the comparison methods, we have plotted in Figure 1 the distributions of the PSNR and SSIM obtained with the optimal parameters on the validation dataset. In Figure 1 all possible comparisons between the proposed methods and the comparison methods are displayed; in particular, in blue and green we plot the PSNR distributions whereas in ocher and violet the SSIM distributions. We observe from the PSNR and SSIM distributions that the validation dataset contains phantoms on which all methods perform poorly (cf. for example Fig. 1(c)); because the validation dataset contains 30 phantoms in total, outliers may explain a higher variation of the PSNR values for our methods in Table 1. If we consider trimmed mean and variances, i.e., excluding the lowest and highest 5% results, we obtain that the trimmed PSNR values on the validation dataset are $31.3\pm 4.24$ for the ZeroShot-PnP algorithm and $30.92\pm 4.15$ for the ZeroShot-$\ell^{1}$-PnP algorithm. The trimmed values have a standard deviation in line with the comparison methods, which are $25.22\pm 3.80$ for the Tikhonov method, $26.78\pm 3.69$ for the ART method, $30.11\pm 2.40$ for the DIP and $28.21\pm 3.14$ for the PP-MPI. Finally, we observe that the independent validation on the hybrid dataset of $\lambda_{\mathrm{Tik}}$ and of the parameters $\mu_{0}$ for the ZeroShot-$\ell^{1}$-PnP algorithms yields parameters $\lambda_{\mathrm{Tik}}=5\cdot 10^{5}$ and $\mu_{0}=2\cdot 10^{5}$ (resp. $\mu_{0}=3\cdot 10^{5}$) of the same order of magnitude and that are not too far away; this is interesting in view of the observation in Equation (15) where we noted that $\mu_{0}$ is interpretable as $\lambda_{\mathrm{Tik}}$, upon zero initialization of $u_{2}^{0}=u_{3}^{0}=0$.
(a) CAD reference.
(b) Tikhonov.
(c) ART.
(d) DIP.
(e) PP-MPI.
(f) ZeroShot-PnP.
(g) ZeroShot-$\ell^{1}$-PnP.
@@
(a) CAD reference.
(b) Tikhonov.
(c) ART.
(d) DIP.
(e) PP-MPI.
(f) ZeroShot-PnP .
(g) ZeroShot-$\ell^{1}$-PnP.
@@
(a) CAD reference.
(b) Tikhonov.
(c) ART.
(d) DIP.
(e) PP-MPI.
(f) ZeroShot-PnP.
(g) ZeroShot-$\ell^{1}$-PnP.
@@
Results.
We apply the proposed methods and the comparison methods to the Open MPI Dataset. The reconstructions where performed with the parameters validated on the hybrid dataset (Table 1). Selected slices of the reconstructions of the three phantoms are displayed in Figures 23 and 4.
@@ Results.
Tikhonov ART DIP PP-MPI [4] with $\ell^{1}$-Prior without $\ell^{1}$-Prior
Shape $\mathrm{PSNR}_{\mathrm{max}}$ 22.64 19.73 28.96 17.65 31.84 31.87
Phatom $\mathrm{SSIM}_{\mathrm{max}}$ 0.489 0.679 0.871 0.654 0.954 0.954
Resolution $\mathrm{PSNR}_{\mathrm{max}}$ 30.64 31.80 32.89 30.69 32.1 32.09
Phantom $\mathrm{SSIM}_{\mathrm{max}}$ 0.593 0.695 0.737 0.658 0.731 0.730
Concentr. $\mathrm{PSNR}_{\mathrm{max}}$ 36.14 39.47 37.78 35.32 37.54 37.45
Phatom $\mathrm{SSIM}_{\mathrm{max}}$ 0.518 0.471 0.441 0.522 0.578 0.577
Table 2: Results obtained with the ZeroShot-$\ell^{1}$-PnP, the Tikhonov, the ART, the DIP and the PP-MPI methods on the 3D Open MPI dataset. We display the $\mathrm{PSNR}_{\mathrm{max}}$ and $\mathrm{SSIM}_{\mathrm{max}}$ scores for the reconstructions obtained using the validated parameters in Table 1.
An overview of the results obtained in terms of $\mathrm{PSNR}_{\mathrm{max}}$ and $\mathrm{SSIM}_{\mathrm{max}}$ are displayed in Tab. 2. We can observe that the ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP methods achieve similar results, and in some instances the additional $\ell^{1}$-prior can increase the PSNR score (cf. shape phantom). In general, the additional $\ell^{1}$ prior does not appear to play a big role in the reconstruction, suggesting that for many applications, the ZeroShot-PnP algorithm could be sufficient. The reconstruction quality according to SSIM of ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP are higher than the baseline Tikhonov, the ART method and the previous PP-MPI approach. The same holds true for the PSNR metric except for the concentration phantom, for which the ART method obtained the highest PSNR score (cf. Table 2). Concerning comparison with the DIP method, we observe from Table 2 that the proposed methods achieve higher PSNR values on the shape phantom, whereas it achieves lower PSNR values on the resolution and concentration phantoms; the SSIM score achieved by the proposed methods are higher also on the concentration phantoms.
However, the DIP takes considerably more time to perform a reconstruction (cf. runtime in Table 3) and involves a random initial state $z$ on which the quality of the reconstruction depends. Moreover, comparing the DIP reconstruction in Figure 4(d) with the CAD reference 4(a) and the ZeroShot-$\ell^{1}$-PnP reconstruction 4(g), it is possible that even if the PSNR value is higher than the value obtained with the proposed methods, some features of the phantom may not be reconstructed by the DIP.
@@
Because the $19\times 19\times 19$ resolution of the reconstructions, the visual quality assessment of the differences between reconstructions can be difficult. For a better understanding of the differences we have plotted in Figure 5 a 1D line plot that compares the graph of the reconstructions and the reference CAD model. In Figure 5 we can see that the Tikhonov method smooths out the reference and takes negative values, whereas the ART, the DIP and the PP-MPI tend to overshoot when compared to the ZeroShot-PnP reconstruction. The ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP lines are almost completely overlapping. We observe that the DIP can present artifacts in the reconstruction (cf. Figure 4(d) and 4(a)) and that the ART method can achieve good results in some phantoms (e.g., the PSNR measure for the concentration phantom in Table 2).
All 2D slices for each of the reconstructed phantom using the proposed and the comparison methods can be found in the Supplementary Material.
@@
Table 3: Wall clock runtime table of the examined methods. The times are computed by averaging 100 reconstructions on the Open MPI dataset. We observe the proposed ZeroShot-PnP and ZeroShot-$\ell^{1}$-PnP take the least amount of time for the reconstructions excluding the Tikhonov method.
Runtimes.
@@
3.4 Experiment 2: Study on the Differences with the PP-MPI Approach.
In this experiment we study more closely the differences between the previous PP-MPI approach and the ZeroShot-$\ell^{1}$-PnP approaches proposed for MPI in this paper. The main difference between the PP-MPI approach and the ZeroShot-$\ell^{1}$-PnP lies in the denoiser employed; the employment of the $\ell^{1}$-HQS-splitting in Equations (9)-(11) is directly related to the choice of the ZeroShot-Denoiser [77]. Indeed, the ZeroShot-Denoiser can take as input the noise level map of the reconstruction. To show the benefit of using the ZeroShot-$\ell^{1}$-PnP algorithm in combination with the ZeroShot-Denoiser (as proposed), we consider in this experiment the following additional method: we substitute the PP-MPI-Denoiser [4] in the $\ell^{1}$-HQS-splitting in place of the ZeroShot-Denoiser.
Preprocessing.
@@
Parameter Validation.
We choose the parameters according to Section 2.4. The validated starting parameters $\mu_{0}$ as well as the resulting number of iterations are displayed in Table 4. We observe from Table 4 that the proposed ZeroShot-PnP and ZeroShot-$\ell^{1}$-PnP achieve the highest PSNR and SSIM scores on the validation dataset. Moreover, we observe that, on the validation dataset, the HQS-splittings (with and without $\ell^{1}$-prior) combined with the PP-MPI-Denoiser have lower PSNR scores when compared to the original PP-MPI method and the proposed methods; nevertheless, the $\ell^{1}$-HQS-splitting with the PP-MPI-Denoiser has a lower number of validated iterations $n_{\mathrm{it}}$ and higher SSIM score when compared with the PP-MPI method.
Hybrid Dataset:
PP-MPI-Denoiser [4] Proposed Denoiser [77]
PP-MPI HQS-Splitting Eq. (12) $\ell^{1}$-HQS-Splitting Eq. (9)-(11) ZeroShot-PnP ZeroShot-$\ell^{1}$-PnP
Param. $\mathrm{SNR}_{\mathrm{db}}=$22 $\mu_{0}=5\cdot 10^{11}$ $\mu_{0}=8\cdot 10^{11}$ $\mu_{0}=2\cdot 10^{5}$ $\mu_{0}=3\cdot 10^{5}$
Iteration $n_{\mathrm{it}}=$1607 $n_{\mathrm{it}}8$ $n_{\mathrm{it}}=7$ $n_{\mathrm{it}}=7$ $n_{\mathrm{it}}=5$
PSNR $28.13\pm 3.75$ $27.06\pm 4.80$ $26.92\pm 4.78$ $30.45\pm 5.04$ $29.77\pm 5.82$
SSIM $0.408\pm 0.217$ $0.483\pm 0.208$ $0.493\pm 0.213$ $0.788\pm 0.141$ $0.766\pm 0.189$
Open MPI Dataset:
Phantom Metric PP-MPI-Denoiser [4] Proposed Denoiser [77]
PP-MPI [4] HQS-Spl. Eq. (12) $\ell^{1}$-HQS-Spl. Eq. (9)-(11) ZeroShot-PnP ZeroShot-$\ell^{1}$-PnP
Shape $\mathrm{PSNR}_{\mathrm{max}}$ 17.65 26.91 27.45 31.87 31.84
Phatom $\mathrm{SSIM}_{\mathrm{max}}$ 0.654 0.712 0.768 0.954 0.954
Resolution $\mathrm{PSNR}_{\mathrm{max}}$ 30.69 27.46 26.44 32.09 32.1
Phantom $\mathrm{SSIM}_{\mathrm{max}}$ 0.658 0.0003 0.0574 0.730 0.731
Concentr. $\mathrm{PSNR}_{\mathrm{max}}$ 35.32 33.43 32.86 37.45 37.54
Phatom $\mathrm{SSIM}_{\mathrm{max}}$ 0.522 0.0041 0.0543 0.577 0.578
Table 4: Validation on the hybrid dataset in Section 2.5 and qualitative comparison on the Open MPI Dataset of the PP-MPI method, the proposed algorithms and the joint $\ell^{1}$-HQS splitting scheme with the PP-MPI-Denoiser. We observe that the proposed ZeroShot-PnP and ZeroShot-$\ell^{1}$-PnP algorithms yields the higher PSNR and SSIM score both on the hybrid dataset and on the Open MPI Dataset.
Results.
We have applied the algorithms with the validated parameters on the Open MPI Dataset. From the results in Table 4, we observe that the proposed ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP algorithms achieve the highest PSNR and SSIM scores. On the validation dataset (cf. Table 4) we have observed that employing the PP-MPI-Denoiser in combination with the $\ell^{1}$-HQS-splitting scheme leads to a higher SSIM score. This results is only partially achieved on the real Open MPI Dataset: from Table 4 it is visible that the $\ell^{1}$-HQS-splitting with the PP-MPI-Denoiser improves the SSIM of the shape phantom when compared with the PP-MPI method, but at the same time, it fails in reconstructing the resolution and concentration phantoms (the SSIM scores are low). In particular, this suggests that the proposed $\ell^{1}$-HQS-splitting scheme works well in combination with the ZeroShot-Denoiser, and the automatic parameter update proposed.
3.5 Experiment 3: Reconstructions with Decreasing Levels of Preprocessing for the ZeroShot-$\ell^{1}$-PnP.
@@
ZeroShot PSNR $30.45\pm 5.04$ $29.73\pm 4.97$ $33.90\pm 6.24$
PnP SSIM $0.788\pm 0.141$ $0.805\pm 0.091$ $0.831\pm 0.124$
$n_{\mathrm{it}}$ 7 8 9
$\mu_{0}$ $2\cdot 10^{5}$ $7\cdot 10^{8}$ $10^{9}$
ZeroShot PSNR $29.77\pm 5.82$ $29.06\pm 5.83$ $32.40\pm 5.01$
$\ell^{1}$-PnP SSIM $0.766\pm 0.189$ $0.802\pm 0.164$ $0.810\pm 0.161$
$n_{\mathrm{it}}$ 5 4 75
@@
Table 5: The ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP algorithms validated with different levels of preprocessing.
Phantom Method Quantity Pre-Processing
Whitening No Whitening No Whitening
rSVD K=2000 rSVD K=2000 rSVD K=6859
Shape ZeroShot $\mathrm{PSNR}_{\mathrm{max}}$ 31.84 31.10 31.30
Phantom PnP $\mathrm{SSIM}_{\mathrm{max}}$ 0.954 0.906 0.950
ZeroShot $\mathrm{PSNR}_{\mathrm{max}}$ 31.87 31.14 29.8
$\ell^{1}$-PnP $\mathrm{SSIM}_{\mathrm{max}}$ 0.954 0.947 0.956
Resolution ZeroShot $\mathrm{PSNR}_{\mathrm{max}}$ 32.10 32.08 31.91
Phantom PnP $\mathrm{SSIM}_{\mathrm{max}}$ 0.731 0.275 0.7132
ZeroShot $\mathrm{PSNR}_{\mathrm{max}}$ 32.09 32.09 31.04
$\ell^{1}$-PnP $\mathrm{SSIM}_{\mathrm{max}}$ 0.730 0.752 0.632
Concentration ZeroShot $\mathrm{PSNR}_{\mathrm{max}}$ 37.54 37.22 39.32
Phantom PnP $\mathrm{SSIM}_{\mathrm{max}}$ 0.578 0.308 0.546
ZeroShot $\mathrm{PSNR}_{\mathrm{max}}$ 37.45 37.16 38.78
$\ell^{1}$-PnP $\mathrm{SSIM}_{\mathrm{max}}$ 0.577 0.568 0.489
Table 6: The ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP reconstruction on the 3D Open MPI Dataset with the validated parameters in Table 5 and metric computed with different levels of preprocessing.
Results.